The electric field in a region is radially outward with magnitude E= At. Find the charge contains in a sphere of radius 20 Cm .given A= 100v/m2

Grade:12

1 Answers

Vikas TU
6 years ago
 Electric Field = 100 r v/m2
            Radius = 20Cm
Electric Field by A sphere = kq/r2
  • 100 r = kq/r2
  • 100r3= kq
  • 100 x 8000 = 9 x 109 x q
  • q = (8x 105)/( = 9 x 109 )
  • q = 8.8 x 10-5